OpenStudy (anonymous):

Simplify: (-a^2b div 2b^-1)^3

OpenStudy (anonymous):

i can answer ur question

OpenStudy (anonymous):

reply something so that i can continue

OpenStudy (anonymous):

((-a )^2 * b / 2b^-1 ))^3 \[-a^{2}timesb\] \[\div (2b)^{-1}\] whole tp the power of 3

OpenStudy (anonymous):

as we know -a^2 = -a * -a = a^2 now, 2b^-1 = 1/2b so now e get a^2 * b/ 1/2b a^2 * b * 2b/1 (2(ab)^2 )^3 8(ab)^6 is the answer if there is any mistake pardon me for that

OpenStudy (anonymous):

have u got the answer correct?

OpenStudy (anonymous):

i wil show u like this